# Do Large Language Models Have Compositional Ability? An Investigation into Limitations and Scalability

### v1

@@ Abstract
Large language models (LLMs) have emerged as powerful tools for many AI problems and exhibit remarkable in-context learning (ICL) capabilities. Compositional ability, solving unseen complex tasks that combine two or more simple tasks, is an essential reasoning ability for Artificial General Intelligence. Despite LLM’s tremendous success, how they approach composite tasks, especially those not encountered during the pretraining phase, remains an open question and largely ununderstood. In this study, we delve into the ICL capabilities of LLMs on composite tasks, with only simple tasks as in-context examples. We develop a test suite of composite tasks that include linguistic and logical challenges and perform empirical studies across different LLM families. We observe that models exhibit divergent behaviors: (1) For simpler composite tasks that apply distinct mapping mechanisms to different input segments, the models demonstrate decent compositional ability, while scaling up the model enhances this ability; (2) for more complex composite tasks that involving reasoning multiple steps, where each step represent one task, models typically underperform, and scaling up generally provide no improvements. We offer theoretical analysis in a simplified setting, explaining that models exhibit compositional capability when the task handles different input parts separately. We believe our work sheds new light on the capabilities of LLMs in solving composite tasks regarding the nature of the tasks and model scale. Our dataset and code are available at https://github.com/OliverXUZY/LLM_Compose.
@@ 1 Introduction
In recent years, large language models (LLM) have revolutionized the natural language processing (NLP) and general AI community. Recent advances, including ChatGPT (OpenAI, 2022), GPT4 (OpenAI, 2023), and Claude 3 (Anthropic, 2024) have shown success in various fields. As model scale increases, larger models exhibit new behavior known as emergence ability. One remarkable emergence is the in-context learning ability (ICL) (Brown et al., 2020), where a model can solve new tasks given only a few examples as input, without any parameters update. However, despite recent success, how LLM solves complex reasoning tasks, particularly not seen in pre-training, remains an open question and largely lacks understanding.
In this paper, we focus on the problem of how LLMs tackle composite tasks that incorporate multiple simple tasks. Specifically, we investigate whether a model trained/in-context learned on individual tasks can effectively integrate these skills to tackle combined challenges, which are intuitive and simple for humans. For instance, in Figure 1, if a human is given examples where words following an asterisk (*) will be capitalized and words surrounded by parenthesis will be permuted, one can also know words following an asterisk (*) surrounded by parenthesis will be capitalized and permuted simultaneously. This basic generalization seems trivial, yet we observe that LLMs fail to generalize this way.
Figure 1: Inconsistent performance in GPT-4. Consider two simple tasks: If a word is followed by an asterisk (*), capitalize the letter. If two words are surrounded by parentheses, swap the positions. GPT-4 correctly solves two simple tasks based on demonstrations (left). The composite tasks have test input with both asterisk (*) and parenthesis. The correct answer should be output: SPORTS PIE. However, GPT-4 fails to solve the composite tasks (right). The same failure was observed in Claude 3.
Compositional ability is an active problem in the AI community and is crucial for advancing Artificial General Intelligence (AGI). Recent studies have made significant contributions to the understanding of this area. Dziri et al. (2023) formulate compositional tasks as computation graphs to quantify each task’s complexity level. Power et al. (2022) show that models may develop generalization capabilities when trained extensively, beyond the point of overfitting, highlighting a phenomenon known as “grokking”. An et al. (2023b) examines how LLMs acquire abstract reasoning and achieve compositional generalization in a linguistic context through ICL by testing LLMs on tasks that involve translating a formal language with a custom grammar. Although these studies offer insight, how LLMs compose tasks together is still not fully understood, especially in the ICL setting. Moreover, the absence of a solid theoretical framework in these discussions needs to be investigated in relation to the underlying mechanisms of such behaviors.
@@ 1.1 Related Work
LLMs are often Transformer-based (Vaswani et al., 2017) equipped with the enormous size of parameters and pretrained on vast training data. One key property that makes such LLM successful is called scaling law: Increasing the scale of language models (pretraining data scale, model parameters) can lead to better performance in downstream tasks. One ability that emerges as the model scale increases is in-context learning (ICL) (Brown et al., 2020). Given a sequence of labeled examples and a testing example (combined as a prompt), the model can construct new predictors for testing examples without further parameter updates Dong et al. (2022). Scaling law was first proposed by Kaplan et al. (2020) and then followed up by Hoffmann et al. (2022), emphasizing both the scale of models and training data. Recent works show LLMs with larger scales have distinct behavior compared to smaller language models (Wei et al., 2023b; Shi et al., 2023b). This work investigates experiments and analyses how LLM can exhibit compositional ability in ICL.
@@ Appendix A Related Work
Large language model. LLMs are often Transformer-based (Vaswani et al., 2017) equipped with the enormous size of parameters and pretrained on vast training data. Typical LLMs includes BERT  Devlin et al. (2019), PaLM Chowdhery et al. (2022), LLaMATouvron et al. (2023), ChatGPT (OpenAI, 2022), GPT4 (OpenAI, 2023). Pretraining methods include masked language modeling (Devlin et al., 2019; Liu et al., 2019), contrastive learning (Gao et al., 2021; Shi et al., 2023a; Sun et al., 2023; 2024) and auto-regressive pretraining (Radford et al., 2018; 2019). Several works (Madasu & Srivastava, 2022; Alajrami et al., 2023) investigate the effects of pretraining on language models. Adapting LLMs to various downstream tasks has received significant attention, e.g., adaptor Hu et al. (2022; 2023); Zhang et al. (2023a); Luo et al. (2024), prompt tuning (Lester et al., 2021; Li & Liang, 2021; Wei et al., 2023a; Gu et al., 2024c), multitask finetuning Sanh et al. (2022); Wang et al. (2023b); Xu et al. (2023; 2024b), instruction tuning Chung et al. (2022); Mishra et al. (2022), in-context learning (Min et al., 2022b; Dong et al., 2022; Yao et al., 2023), low-rank adaptation Hu et al. (2022); Zeng & Lee (2024); Hu et al. (2024), reinforcement learning from human feedback (RLHF) Ouyang et al. (2022) and inference acceleration Gu et al. (2024b; d).
@@
Emergence of compositional ability. Scaling law was first proposed by Kaplan et al. (2020) and then followed up by Hoffmann et al. (2022), emphasizing both the scale of models and training data. Sometimes, increasing scale can lead to new behaviors of LLMs, termed emergent abilities (Wei et al., 2022; Arora & Goyal, 2023), such as domain generalization Shi et al. (2024), math reasoning Gu et al. (2024a), spatial reasoning Wang et al. (2024) and so on. Recent works show LLMs with larger scales have distinct behavior compared to smaller language models (Wei et al., 2023b; Shi et al., 2023b). These behaviors can have positive or negative effects on performance. Solving complex tasks and reasoning is an active problem in the AI community Huang & Chang (2022). There is a line of empirical works investigating the compositional ability in linguistic fashion (Kim & Linzen, 2020; Levy et al., 2022; An et al., 2023a; b; Xu et al., 2024a). LLMs are capable of learning abstract reasoning (e.g., grammar) to perform new tasks when finetuned or given suitable in-context examples. In our work, we include linguistic experiments as part of our testing suite, illustrating LLMs’ compositional ability. Ye et al. (2023); Berglund et al. (2023); Dziri et al. (2023) show LLMs will have difficulties solving tasks that require reasoning. Berglund et al. (2023) studies that LLMs trained on “A is B” fail to learn “B is A”. In our work, we conduct similar experiments showing LLMs will fail on composite if different steps of logical rules are mixed.

### v2

@@ Abstract
Large language models (LLMs) have emerged as powerful tools for many AI problems and exhibit remarkable in-context learning (ICL) capabilities. Compositional ability, solving unseen complex tasks that combine two or more simple tasks, is an essential reasoning ability for Artificial General Intelligence. Despite the tremendous success of LLMs, how they approach composite tasks, especially those not encountered during the pretraining phase, remains an open and largely underexplored question. In this study, we delve into the ICL capabilities of LLMs on composite tasks, with only simple tasks as in-context examples. We develop a test suite of composite tasks including linguistic and logical challenges and perform empirical studies across different LLM families. We observe that models exhibit divergent behaviors: (1) For simpler composite tasks that apply distinct mapping mechanisms to different input segments, the models demonstrate decent compositional ability, while scaling up the model enhances this ability; (2) for more complex composite tasks involving reasoning multiple steps, where each step represents one task, models typically underperform, and scaling up generally provides no improvements. We offer theoretical analysis in a simplified setting, explaining that models exhibit compositional capability when the task handles different input parts separately. We believe our work sheds new light on the capabilities of LLMs in solving composite tasks regarding the nature of the tasks and model scale. Our dataset and code are available at https://github.com/OliverXUZY/LLM_Compose.
@@ 1 Introduction
In recent years, large language models (LLMs) have revolutionized the natural language processing (NLP) and general AI community. Recent advances, including ChatGPT (OpenAI, 2022), GPT4 (OpenAI, 2023), and Claude 3 (Anthropic, 2024) have shown success in various fields. As model scale increases, larger models exhibit new behavior known as emergence ability. One remarkable emergence is the in-context learning ability (ICL) (Brown et al., 2020), where a model can solve new tasks given only a few examples as input, without any parameter updates. However, despite recent success, how LLMs solve complex reasoning tasks, particularly not seen in pre-training, remains an open question and largely lacks understanding.
In this paper, we focus on the problem of how LLMs tackle composite tasks that incorporate multiple simple tasks. Specifically, we investigate whether a model trained and in-context learned on individual tasks can effectively integrate these skills to tackle combined challenges, which are intuitive and simple for humans. For instance, in Figure 1, if a human is given examples where words following an asterisk (*) will be capitalized and words surrounded by parenthesis will be permuted, one can also understand words following an asterisk (*) surrounded by parenthesis will be capitalized and permuted simultaneously. This basic generalization seems trivial, yet we observe that LLMs fail to generalize in this way.
Figure 1: Inconsistent performance in GPT-4. Consider two simple tasks: If a word is followed by an asterisk (*), capitalize the letter. If two words are surrounded by parentheses, swap the positions. GPT-4 correctly solves two simple tasks based on demonstrations (left). The composite tasks have test inputs with both asterisk (*) and parenthesis. The correct answer should be output: SPORTS PIE. However, GPT-4 fails to solve the composite tasks (right). The same failure was observed in Claude 3.
Compositional ability is an active problem in the AI community and is crucial for advancing Artificial General Intelligence (AGI). Recent studies have made significant contributions to the understanding of this area. Dziri et al. (2023) formulate compositional tasks as computation graphs to quantify each task’s complexity level. Power et al. (2022) show that models may develop generalization capabilities when trained extensively, beyond the point of overfitting, highlighting a phenomenon known as “grokking”. An et al. (2023b) examines how LLMs acquire abstract reasoning and achieve compositional generalization in a linguistic context through ICL by testing LLMs on tasks that involve translating a formal language with a custom grammar. Although these studies offer insight, how LLMs compose tasks together is still not fully understood, especially in the ICL setting. Moreover, the absence of a solid theoretical framework in these discussions needs to be investigated concerning the underlying mechanisms of such behaviors.
@@ 1.1 Related Work
LLMs are often Transformer-based (Vaswani et al., 2017) equipped with the enormous size of parameters and pretrained on vast training data. One key property that makes such LLM successful is called scaling law: Increasing the scale of language models (pretraining data scale, model parameters) can lead to better performance in downstream tasks. One ability that emerges as the model scale increases is in-context learning (ICL) (Brown et al., 2020). Given a sequence of labeled examples and a testing example (combined as a prompt), the model can construct new predictors for testing examples without further parameter updates Dong et al. (2022). Scaling law was first proposed by Kaplan et al. (2020) and then followed up by Hoffmann et al. (2022), emphasizing both the scale of models and training data. Recent works show LLMs with larger scales have distinct behaviors compared to smaller language models (Wei et al., 2023b; Shi et al., 2023b). This work investigates experiments and analyses how LLM can exhibit compositional ability in ICL.
@@ Appendix A Related Work
Large language model. LLMs are often Transformer-based (Vaswani et al., 2017) equipped with the enormous size of parameters and pretrained on vast training data. Typical LLMs includes BERT  Devlin et al. (2019), PaLM Chowdhery et al. (2022), LLaMATouvron et al. (2023), ChatGPT (OpenAI, 2022), GPT4 (OpenAI, 2023). Pretraining methods include masked language modeling (Devlin et al., 2019; Liu et al., 2019), contrastive learning (Gao et al., 2021; Shi et al., 2023a; Sun et al., 2023; 2024) and auto-regressive pretraining (Radford et al., 2018; 2019). Several works (Madasu & Srivastava, 2022; Alajrami et al., 2023) investigate the effects of pretraining on language models. Adapting LLMs to various downstream tasks has received significant attention, e.g., adaptor (Hu et al., 2022; 2023; Zhang et al., 2023a; Luo et al., 2024), prompt tuning (Lester et al., 2021; Li & Liang, 2021; Wei et al., 2023a; Gu et al., 2024c), multitask finetuning (Sanh et al., 2022; Wang et al., 2023b; Xu et al., 2023; 2024c), instruction tuning Chung et al. (2022); Mishra et al. (2022), in-context learning (Min et al., 2022b; Dong et al., 2022; Yao et al., 2023), low-rank adaptation Hu et al. (2022); Zeng & Lee (2024); Hu et al. (2024), reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022) and inference acceleration (Gu et al., 2024b; d; Xu et al., 2024a).
@@
Emergence of compositional ability. Scaling law was first proposed by Kaplan et al. (2020) and then followed up by Hoffmann et al. (2022), emphasizing both the scale of models and training data. Sometimes, increasing scale can lead to new behaviors of LLMs, termed emergent abilities (Wei et al., 2022; Arora & Goyal, 2023), such as domain generalization Shi et al. (2024a), math reasoning Gu et al. (2024a), spatial reasoning Wang et al. (2024) and so on. Recent works show LLMs with larger scales have distinct behavior compared to smaller language models (Wei et al., 2023b; Shi et al., 2023b; 2024b). These behaviors can have positive or negative effects on performance. Solving complex tasks and reasoning is an active problem in the AI community Huang & Chang (2022). There is a line of empirical works investigating the compositional ability in linguistic fashion (Kim & Linzen, 2020; Levy et al., 2022; An et al., 2023a; b; Xu et al., 2024b). LLMs are capable of learning abstract reasoning (e.g., grammar) to perform new tasks when finetuned or given suitable in-context examples. In our work, we include linguistic experiments as part of our testing suite, illustrating LLMs’ compositional ability. Ye et al. (2023); Berglund et al. (2023); Dziri et al. (2023) show LLMs will have difficulties solving tasks that require reasoning. Berglund et al. (2023) studies that LLMs trained on “A is B” fail to learn “B is A”. In our work, we conduct similar experiments showing LLMs will fail on composite if different steps of logical rules are mixed.